# You Need to Pay Better Attention: Rethinking the Mathematics of Attention Mechanism

### v2

@@ Abstract
Scaled Dot Product Attention (SDPA) is the backbone of many modern deep-learning models. It is so versatile that it has been used in natural language, vision, and multi-modal domains with very little change compared to its original formulation. This paper discusses why the current formulation is inefficient by delving into the mathematical details of the attention mechanism. We propose three improvements to mitigate these inefficiencies, thereby, introducing three enhanced attention mechanisms: Optimised, Efficient, and Super Attention. Optimised and Efficient Attention have one and two matrix multiplications fewer per head, respectively, and 25% and 50% fewer parameters, respectively, than standard SDPA, but perform similarly to standard SDPA in both vision and natural language tasks. They can be used in all applications where SDPA is used while offering smaller model sizes and faster training and inference without noticeable loss in performance. Super Attention introduces a new linear transformation on the values, transforming them from the left. It outperforms standard SPDA on vision and natural language tasks by up to 17% while having one fewer matrix multiplication per head and 25% fewer parameters than standard SDPA. Consequently, it is also faster than standard SDPA. Super Attention is ideal in applications where the attention layer’s context length is fixed, such as Vision Transformers. In addition to providing mathematical reasoning, we evaluate the presented attention mechanisms on several datasets including MNIST, CIFAR100, ImageNet, IMDB Movie Reviews, and Amazon Reviews datasets, as well as combined Europarl and Anki English-Spanish datasets for neural machine translation.
@@ 1 Introduction
Not many ideas have had as profound an effect on the field of Artificial Intelligence (AI) as the attention mechanism (Bahdanau et al., 2015). Introduced as a method to improve machine translation, the attention mechanism revolutionised the way neural networks process and interpret data. By allowing models to focus on specific parts of the input while disregarding irrelevant information, it mimics a form of cognitive attention in humans. It not only enhanced the capability and efficiency of Language Models (LM) but also paved the way for the development of advanced AI architectures like the Transformer model (Vaswani et al., 2017).
@@
Figure 1: Standard multi-head scaled dot product attention (1(a)) alongside the four proposed modifications: Optimised Attention (1(b)), Efficient Attention (1(c)), and Super Attention (1(d)). The “Linear” block denotes a linear transformation right while “Linear*” denotes a linear transformation from left.
These advances have had far-reaching impacts, extending beyond Natural Language Processing (NLP) to other areas such as image recognition (Dosovitskiy et al., 2021), autonomous systems (Mott et al., 2019), healthcare (Choi et al., 2016), and multi-modal application Xu et al. (2023).
The formulation of SDPA in all these domains has undergone very little change compared to the original formulation of Vaswani et al. (2017). Instead, “The bigger the better" has been the prevailing maxim in AI in the last few years. Larger Language Models (LLM), such as Llama 3 (Touvron et al., 2023a, b), GPT-4 (Achiam et al., 2023), and Gemini (Anil et al., 2023) have demonstrated unprecedented capabilities in multi-modal domains.
The behemothic sizes of these models have introduced numerous challenges, such as expensive and slow training and inference, leading to secondary problems such as high carbon emission (Dhar, 2020). Furthermore, such models are impossible not only to run but even to store on edge devices such as smartphones, consumer laptops, and even powerful personal workstations.
In recent years, there have been numerous attempts to address this problem using post-training techniques, like quantisation (Jacob et al., 2018), Low-Rank Adaptation (LoRA) (Hu et al., 2022), Quantised LoRA (QLoRA) (Dettmers et al., 2023), and sparsification (Ashkboos et al., 2024). There have been also attempts to optimise the speed and GPU utilisation of attention-based models. Notable examples include Flash Attention Dao et al. (2022) and its successor, Flash Attention 2 Dao (2024).
All these approaches focus on techniques to improve the performance of attention-based models without altering the attention mechanism. In this paper, we look into the attention mechanism itself and put forward three attention mechanisms, Optimised Attention, Efficient Attention, and Super Attention. Our contributions are founded on two intuitive principles:
  1. 1.
    Two consecutive linear transformations result in a linear transformation.
  2. 2.
    A linear kernel between each input increases learning.
Using Principle 1, Optimised Attention omits the values’ linear transformation, (see LABEL:subfig:_OptimisedAttention and Eq. 4) while preserving the learning capabilities. We use Principle 1 once more in Efficient Attention to not only omit the values’ linear transformation but also that of the keys (see Fig. 1(c) and Eq. 5). Finally, Super Attention uses Principle 2 and introduces a new learnable kernel $W^{A}$ in between the attention scores and values (see Fig. 1(d) and Eq. 22), thus enhancing the performance of SDPA in both vision and NLP tasks compared to standard attention, while being more efficient and having fewer parameters. We validate our proposed algorithms on image classification on MNIST, CIFAR100, and ImageNet datasets, natural language sentiment classification on IMDB and Amazon Reviews datasets, and Neural Machine Translation (NMT) on the combined Europarl and Anki English-to-Spanish translation dataset.
Contribution.
In summary, we introduce:
@@
    Optimised Attention in LABEL:subsec:Merging_WV_and_WO, which reduces the attention layer’s size by 1/4 and its computational cost by $h$ matrix multiplication, where $h$ is the number of heads; thus, reducing its training and inference time by 2.5–7.5%, while performing similarly to standard attention in terms of learning performance.
@@
    Efficient Attention in LABEL:subsec:Merging_WQ_and_WK, which is our most efficient attention mechanism. It reduces the attention layer’s size by 1/2 and its computational cost by $2h$ matrix multiplications; thus, reducing its training and inference time by 5–15%, while performing similarly to standard attention in terms of learning performance.
@@
    Super Attention in LABEL:subsec:Introducing_WA, which is our most capable attention mechanism, outperforming standard attention by up to 2–17% in vision and NLP. It reduces the attention layer’s size by 1/4 and its computational cost by $h$ matrix multiplications, when the context length is equal to or smaller than the model dimension, thereby reducing the training and inference time by 2.5–10%.
@@ 2 Preliminaries
Before defining the SDPA, let us introduce the notations used throughout the paper. For natural numbers $d_{m},d_{k}\in\mathbb{N}$, we denote the $d_{m}$-dimensional real vectors space by $\mathbb{R}^{d_{m}}$ and the set of all real $d_{m}\times d_{k}$ matrices by $\mathbb{R}^{d_{m}\times d_{k}}$, noting that all matrices can be regarded as 2D tensors and vice versa. Given a set $\mathcal{A}\subseteq\mathbb{R}^{d_{m}}$, we denote the smallest real vector space containing $\mathcal{A}$ by $\operatorname*{span}(\mathcal{A})$. Similarly, given a matrix $W\in\mathbb{R}^{d_{m}\times d_{k}}$, we denote the smallest real vector space containing the columns of $W$’s by $\operatorname*{span}(W)$. For a subspace $\mathcal{S}\leq\mathbb{R}^{d_{m}}$, the dimension of $\mathcal{S}$, denoted $\dim(\mathcal{S})$, is the size of the largest linearly independent set in $\mathcal{S}$. The rank of a matrix $W\in\mathbb{R}^{d_{m}\times d_{k}}$, denoted $\operatorname{rank}(W)$, is the number of linearly independent columns (or rows) in $W$. The rank-nullity theorem implies that $\operatorname{rank}(W)=\dim(\operatorname*{span}(W))$ and $\operatorname{rank}(W)\leq\min(d_{m},d_{k})$. [^1] For a detailed introduction to these see (Meyer, 2023, Chapters 2 & 4).
We use the definition of the attention mechanism used in the implementations of SDPA in machine learning frameworks, such as Torch, JAX, TensorFlow, and Keras.
Definition 1 (Standard Attention).
The (multi-head) scaled dot-product attention on input tensors $Q,K,V\in\mathbb{R}^{\ell\times d_{m}}$ is defined as
@@
Parameters $d_{m},d_{k},d_{v}$ and $h$ are often chosen so that $d_{k}=d_{v}=d_{m}/h$, and in most recent models, including Transformer models, $Q,K$, and $V$ are set to $X$, a single input tensor; whereby, the attention mechanism is called self-attention.
@@ 3 Revising the Attention Mechanism
In this section, we visit and rethink the mathematical underpinnings of the attention mechanism and present enhancements to the attention mechanism to make it more efficient (in terms of number of parameters and computation cost) and more capable (in terms of attaining higher accuracies and lower losses). In particular, we introduce Optimised Attention in Section 3.1, Efficient Attention in Section 3.2, and Super Attention in Section 3.3. We provide the mathematical motivation behind each of the mechanisms in the corresponding sections and evaluate all of them in Section 4.
3.1 Optimised Attention: Absorbing $W^{V}_{i}$’s into $W^{0}$
We start by optimising operations (1) and (4) of the attention mechanism. We do this by absorbing $W^{V}_{1},W^{V}_{2},\dots,W^{V}_{h}$ into $W^{O}$, thereby reducing the computational cost of the attention layer by $h$ matrix multiplications without significantly affecting the performance as we show in Section 4.
In standard attention, the output $O$ can be written as
$\begin{split}O&=(H_{1}\ H_{2}\ \cdots\ H_{h})W^{O}=(S_{1}VW^{V}_{1}\ S_{2}VW^{V}_{2}\ \cdots\ S_{h}VW^{V}_{h})\begin{pmatrix}W^{O}_{1}\\ W^{O}_{2}\\ \vdots\\ W^{O}_{h}\end{pmatrix}\\ &=S_{1}VW^{V}_{1}W^{O}_{1}+S_{2}VW^{V}_{2}W^{O}_{2}+\dots+S_{h}VW^{V}_{h}W^{O}_{h},\end{split}$ (7)
where $W^{O}_{i}$ is the matrix that contains rows $(i-1)d_{v}+1,\dots,id_{v}$ of $W^{O}$ for $i=1,2,\dots,h$. By the rank-nullity theorem, for each head, we have that
$\begin{split}\dim(\operatorname*{span}(VW^{V}_{i}W^{O}_{i}))&=\operatorname{rank}(VW^{V}_{i}W^{O}_{i})\leq\operatorname{rank}(W^{V}_{i}W^{O}_{i}),\\ &\leq\min(\operatorname{rank}(W^{V}_{i}),\operatorname{rank}(W^{O}_{i}))=\min(d_{m},d_{v})=d_{v}.\end{split}$
In other words, $VW^{V}_{i}W^{O}_{i}$ has at most $d_{v}$ independent columns, and the linear function $V\mapsto VW^{V}_{i}W^{O}_{i}$ maps the columns of $V$ into a $d_{v}$-dimensional subspace of $\mathbb{R}^{d_{m}}$.
Thus, standard attention uses two consecutive matrix multiplication to embed the columns of $V$ into a $d_{v}$-dimensional subspace of $\mathbb{R}^{d_{m}}$, which is inefficient according to Principle 1 (and as validated in Section 4). As shown in Fig. 1(b) and described in LABEL:def:Merging_WV_and_WO, in Optimised attention, we achieve the same effect by one slicing and one matrix multiplication, thereby reducing the the computational cost of attention during training and inference.
In more detail, instead of multiplying $V$ from the right by $W^{V}_{i}$, we first slice $V$ into $V_{1},\dots,V_{h}$, where $V_{i}$ consists of columns $(i-1)d_{v}+1,\dots,id_{v}$ of $V$, and then, instead of computing $S_{i}VW^{V}_{i}W^{O}_{i}$, we compute $S_{i}V_{i}W^{O}_{i}$, which has fewer parameters and matrix multiplications (see Remark 1). As we show in LABEL:sec:Evaluation, Optimised Attention considerably improves the efficiency of the attention layer without affecting the model’s performance.
Definition 2 (Optimised Attention).
Using the notation of Definition 1, Optimised Attention is the attention mechanism defined by the following set of equations:
@@ 3.1 Optimised Attention: Absorbing $W^{V}_{i}$’s into $W^{0}$
Remark 1.
Optimised Attention is more efficient than standard attention in the sense that it has $h$ matrix multiplication and $d^{2}_{m}$ parameters fewer than standard attention.
@@
Compared to Optimised Attention, standard attention has extra $W^{V}_{1},W^{V}_{2},\dots,W^{V}_{h}$, which are multiplied from the right to $V$, amounting to a total of $d_{m}d_{v}h=d_{m}^{2}$ parameters and $h$ matrix multiplications. ∎
@@ 3.2 Efficient Attention: Absorbing $W^{K}$ into $W^{Q}$
In the last section, we discussed how and why we can remove $W^{O}$ to reduce the computational cost of the attention mechanism. Here, we show that we can do the same for $W^{K}$ to further reduce the attention mechanism’s computational cost. When computing the pre-$\operatorname{softmax}$ attention scores for each head, we have that
$\begin{split}\dim(\operatorname*{span}(\frac{QW_{i}^{Q}{W^{K}}^{\intercal}_{i}{K}^{\intercal}}{d_{k}})&=\operatorname{rank}(QW_{i}^{Q}{W^{K}}^{\intercal}_{i}{K}^{\intercal})\leq\operatorname{rank}(W_{i}^{Q}{W^{K}}^{\intercal}_{i}),\\ &\leq\min(\operatorname{rank}(W_{i}^{Q}),\operatorname{rank}(W^{K}_{i}))=\min(d_{m},d_{k})=d_{k}.\end{split}$
Therefore, SDPA suffers from the same problem as before in Section 3.1. Similarly to what we did in Optimised Attention, we address this by replacing the linear transformation of keys by slicing as depicted in LABEL:subfig:Efficient_Attention and described in Definition 3.
Definition 3 (Efficient Attention).
Using the same notation as Definition 2, Efficient Attention is defined via the following equations:
@@
In Efficient Attention, we do not have $W^{K}_{1},W^{K}_{2},\dots,W^{K}_{h}$, which are applied to $K$ from left. Hence, compared to Optimised Attention, we have reduced the number of matrix multiplications by $h$ and parameters by $d_{m}^{2}$. From this and Remark 1, it follows that Efficient Attention has $h+h=2h$ matrix multiplication and $d_{m}^{2}+d_{m}^{2}=2d_{m}^{2}$ parameters less than standard attention. ∎

### v3

# Cost-Effective Attention Mechanisms for Low Resource Settings: Necessity & Sufficiency of Linear Transformations

@@ Abstract
From natural language processing to vision, Scaled Dot Product Attention (SDPA) is the backbone of most modern deep learning applications. Unfortunately, its memory and computational requirements can be prohibitive in low-resource settings. In this paper, we improve its efficiency without sacrificing its versatility. We propose three attention variants where we remove consecutive linear transformations or add a novel one, and evaluate them on a range of standard NLP and vision tasks. Our proposed models are substantially lighter than standard SDPA (and have 25-50% fewer parameters). We show that the performance cost of these changes is negligible relative to size reduction and that in one case (Super Attention) we succeed in outperforming SDPA by up to 10% while improving its speed and reducing its parameters by 25%.
Cost-Effective Attention Mechanisms for Low Resource Settings:
Necessity & Sufficiency of Linear Transformations
Peyman Hosseini1     Mehran Hosseini2     Ignacio Castro1     Matthew Purver1,3 1School of EECS, Queen Mary University of London, London, UK 2Department of Informatics, King’s College London, London, UK 3Department of Knowledge Technologies, Jožef Stefan Institute, Ljubljana, Slovenia {s.hosseini, i.castro, m.purver}@qmul.ac.uk, mehran.hosseini@kcl.ac.uk
@@ 1 Introduction
Few ideas have had as profound an effect on the field of Artificial Intelligence (AI) as the attention mechanism (Bahdanau et al., 2015). Introduced as a method to improve machine translation, the attention mechanism revolutionized the way neural networks process and interpret data. It mimics a form of cognitive attention in humans by allowing models to focus on specific parts of the input while disregarding irrelevant information. This enhanced the capability and efficiency of Language Models (LM) and paved the way for the development of advanced AI architectures like the Transformer model (Vaswani et al., 2017).
@@
Figure 1: Standard multi-head scaled dot product attention (1(a)) alongside the proposed variations: Optimized Attention (1(b)), Efficient Attention (1(c)), and Super Attention (1(d)). The “Linear” block denotes a linear transformation right while “Linear*” denotes a linear transformation from left.
These advances have had far-reaching impacts, extending beyond Natural Language Processing (NLP) to areas such as image recognition (Dosovitskiy et al., 2021), autonomous systems (Mott et al., 2019), healthcare (Choi et al., 2016), and multi-modal application Xu et al. (2023).
The formulation of SDPA in all these domains has undergone very little change compared to the original formulation of Vaswani et al. (2017). Instead, the prevailing maxim has been “the bigger the better", and Large Language Models (LLM), such as Llama 3 (Touvron et al., 2023a, b), GPT-4 (Achiam et al., 2023), and Gemini (Anil et al., 2023) have demonstrated unprecedented capabilities in multi-modal domains.
The behemothic sizes of these models have introduced numerous challenges. Expensive and slow training and inference have resulted in high carbon emissions (Dhar, 2020); and such models are impossible not only to run but even to store on edge devices such as smartphones, consumer laptops, and even powerful personal workstations.
Numerous attempts have been made to address this problem using post-training techniques, like quantization (Jacob et al., 2018), Low-Rank Adaptation (LoRA) (Hu et al., 2022), Quantized LoRA (QLoRA) (Dettmers et al., 2023), and sparsification (Ashkboos et al., 2024). Others have attempted to optimise the speed and GPU utilization of attention-based models, e.g., Flash Attention 1–3 (Dao et al., 2022; Dao, 2024; Shah et al., 2024). However, all these approaches strive to improve the performance of attention-based models but without altering the attention mechanism.
In this paper, we propose a different approach: modifying the attention mechanism itself. We employ two intuitive principles to design our alternative attention mechanism: (1) two consecutive linear transformations do not introduce non-linearity, and (2) a learnable linear kernel between each two inputs of SDPA enhances learning. We leverage these two principles to propose 3 SDPA variants:
@@
    Optimized Attention (LABEL:subsec:Merging_WV_and_WO, Fig. 1(b)), replaces $W^{V}$ linear transformation with a slicing operation (Principle 1), reducing the parameters in the attention layer by 25% and its computational cost by $h$ matrix multiplications, where $h$ is the number of heads. Optimized Attention reduces the inference time by 2.5–7.5%, with little or no performance degradation (§ 4).
@@
    Efficient Attention (LABEL:subsec:Merging_WQ_and_WK, Fig. 1(c)) replaces $W^{V}$ and $W^{K}$ linear transformations by slicing operations (Principle 1). This reduces the parameters in the attention layer by 50% and its computational cost by $2h$ matrix multiplications, where $h$ is the number of heads. Efficient Attention reduces the inference time by 5–15%, with no/little performance degradation (§ 4).
@@
    Super Attention (LABEL:subsec:Introducing_WA, Fig. 1(d)) introduces a new linear operation $W^{A}$ (Principle 2), which transforms the values $V$ from the left. Super Attention can be used on top of standard or Optimized attentions (i.e., without replacing $W^{V}$ and $W^{K}$). For simplicity, we build Super Attention on top of Efficient Attention. Super Attention reduces the attention layer’s size by $\sim 25\%$ (depending on the attention’s context length) and its computational cost by $h$ matrix multiplications. Super Attention outperforms standard attention by 2–10% in NLP and vision tasks and reduces the training and inference time by 2.5–10% (§ 4).
Our evaluation is comprehensive and compares our proposed attention models with SDPA in the self-attention setting in transformers for multiple datasets and for 4 different tasks, including: (1) Natural Language Sentiment Classification on IMDB and Amazon Reviews datasets; (2) Machine Translation (NMT) on the combined Europarl and Anki English-to-Spanish translation dataset; (3) Generative Language Modeling and Natural Language Inference (NLI) using NanoGPT Karpathy (2022) on the OpenWebText dataset; and to show how these architectural changes generalize to transformers for other modalities, we do complementary experiments for (1) image classification on MNIST, CIFAR100, and ImageNet datasets;
@@ 2 Preliminaries
We start by introducing the notation we use throughout the paper. For natural numbers $d_{m},d_{k}\in\mathbb{N}$, we denote the $d_{m}$-dimensional real vectors space by $\mathbb{R}^{d_{m}}$ and the set of all real $d_{m}\times d_{k}$ matrices by $\mathbb{R}^{d_{m}\times d_{k}}$, noting that all matrices can be regarded as 2D tensors and vice versa. Given a set $\mathcal{A}\subseteq\mathbb{R}^{d_{m}}$, we denote the smallest real vector space containing $\mathcal{A}$ by $\operatorname*{span}(\mathcal{A})$. Similarly, given a matrix $W\in\mathbb{R}^{d_{m}\times d_{k}}$, we denote the smallest real vector space containing the columns of $W$’s by $\operatorname*{span}(W)$. For a subspace $\mathcal{S}\leq\mathbb{R}^{d_{m}}$, the dimension of $\mathcal{S}$, denoted $\dim(\mathcal{S})$, is the size of the largest linearly independent set in $\mathcal{S}$. The rank of a matrix $W\in\mathbb{R}^{d_{m}\times d_{k}}$, denoted $\operatorname{rank}(W)$, is the number of linearly independent columns (or rows) in $W$. The rank-nullity theorem implies that $\operatorname{rank}(W)=\dim(\operatorname*{span}(W))$ and $\operatorname{rank}(W)\leq\min(d_{m},d_{k})$. [^1] For details see (Meyer, 2023, Chapters 2 & 4).
We use the widely-adopted definition of SDPA as implemented in SotA open-source models such as Llama-3 and Mistral, and machine learning frameworks like Torch and JAX. For consistency, we use the same notation as (Vaswani et al., 2017).
Definition 2.1 (Standard Attention).
The (multi-head) scaled dot-product attention on input tensors $Q,K,V\in\mathbb{R}^{\ell\times d_{m}}$ is defined as:
@@
Parameters $d_{m},d_{k},d_{v}$ and $h$ are often chosen so that $d_{k}=d_{v}=d_{m}/h$, and in recent models, including SotA Transformer models, $Q,K$, and $V$ are set to $X$, a single input tensor; whereby, the attention mechanism is called self-attention.
@@ 3 Revising the Attention Mechanism
We introduce our three proposed attention variants and discuss the motivation behind each of them.
3.1 Optimized Attention: Absorbing $W^{V}_{i}$’s into $W^{0}$
In standard attention, the output $O$ of the attention layer can be written as
$O$ $=(H_{1}\ \cdots\ H_{h})W^{O}$ (7)
$=(S_{1}VW^{V}_{1}\ \cdots\ S_{h}VW^{V}_{h})\begin{pmatrix}W^{O}_{1}\\ \vdots\\ W^{O}_{h}\end{pmatrix}$
$=S_{1}VW^{V}_{1}W^{O}_{1}+\cdots+S_{h}VW^{V}_{h}W^{O}_{h},$
where $W^{O}_{i}$ is the matrix with rows $(i-1)d_{v}+1,\dots,id_{v}$ of $W^{O}$ for $i=1,2,\dots,h$. By the rank-nullity theorem, for each head, we have that:
$\dim$ $(\operatorname*{span}(VW^{V}_{i}W^{O}_{i}))$
$=\operatorname{rank}(VW^{V}_{i}W^{O}_{i})\leq\operatorname{rank}(W^{V}_{i}W^{O}_{i}),$
$\leq\min(\operatorname{rank}(W^{V}_{i}),\operatorname{rank}(W^{O}_{i}))$
$=\min(d_{m},d_{v})=d_{v}.$
That is, $VW^{V}_{i}W^{O}_{i}$ has at most $d_{v}$ independent columns, and the linear function $V\mapsto VW^{V}_{i}W^{O}_{i}$ maps the columns of $V$ into a $d_{v}$-dimensional subspace of $\mathbb{R}^{d_{m}}$. Thus, standard attention uses two consecutive matrix multiplications to embed the columns of $V$ into a $d_{v}$-dimensional subspace of $\mathbb{R}^{d_{m}}$, which does not align with Principle (1).
To address this, in Optimized Attention, we absorb $W^{V}_{1},W^{V}_{2},\dots,W^{V}_{h}$ into $W^{O}$ in Eqs. (1) and (4), thus reducing the computational cost of the attention layer by $h$ matrix multiplications at a very limited performance cost–which we evaluate in § 4.
Optimized Attention uses one slicing and one linear transformation (see Fig. 1(b) and Def. 3.1), instead of the two consecutive linear transformations (one downscaling and one upscaling). Specifically, instead of multiplying $V$ from the right by $W^{V}_{i}$, we slice $V$ into $V_{1},\dots,V_{h}$, where $V_{i}$ consists of columns $(i-1)d_{v}+1,\dots,id_{v}$ of $V$, and then, instead of computing $S_{i}VW^{V}_{i}W^{O}_{i}$, we compute $S_{i}V_{i}W^{O}_{i}$, which needs fewer parameters and matrix multiplications (see Rem. 3.2 and § 4.3 for theoretical and empirical evaluations, respectively.)
Definition 3.1 (Optimized Attention).
Using the notation of Def. 2.1, Optimized Attention is defined as follows:
@@ 3.1 Optimized Attention: Absorbing $W^{V}_{i}$’s into $W^{0}$
Remark 3.2.
Optimized Attention is more efficient than standard attention, having $h$ matrix multiplication and $d^{2}_{m}$ parameters less than standard attention.
@@
Compared to Optimized Attention, standard attention has extra $W^{V}_{1},W^{V}_{2},\dots,W^{V}_{h}$, which are multiplied from the right to $V$. This amounts to a total of $d_{m}d_{v}h=d_{m}^{2}$ parameters and $h$ matrix multiplications. ∎
@@ 3.2  Efficient Attention: Absorbing $W^{K}$ into $W^{Q}$
In § 3.1, we discussed our motivation behind dropping $W^{V}$. Here, we do the same for $W^{K}$ to further reduce the computational cost of the attention mechanism. Before this, we note that for the pre-$\operatorname{softmax}$ attention scores for each head, we have:
$\dim(\operatorname*{span}(\frac{QW_{i}^{Q}{W^{K}}^{\intercal}_{i}{K}^{\intercal}}{d_{k}})$
$\ =\operatorname{rank}(QW_{i}^{Q}{W^{K}}^{\intercal}_{i}{K}^{\intercal})\leq\operatorname{rank}(W_{i}^{Q}{W^{K}}^{\intercal}_{i}),$
$\ \leq\min(\operatorname{rank}(W_{i}^{Q}),\operatorname{rank}(W^{K}_{i}))$
$\ =\min(d_{m},d_{k})=d_{k}.$
More precisely, here two linear kernels, $W_{i}^{Q}$ and ${W^{K}}^{\intercal}_{i}$, are stacked– this opposes Principle (1). Thus, following the same approach as in Optimized Attention, we merge ${W^{K}}^{\intercal}_{i}$ into $W_{i}^{Q}$ and replace the $W^{K}_{i}$ linear transformation by slicing as depicted in LABEL:subfig:Efficient_Attention and defined in Def. 3.3.
Definition 3.3 (Efficient Attention).
Using the same notation as Def. 3.1, we define Efficient Attention with the following equations:
@@
In Efficient Attention, we do not have $W^{K}_{1},W^{K}_{2},\dots,W^{K}_{h}$, which are applied to $K$ from left. Hence, we reduce the number of matrix multiplications by $h$ and parameters by $d_{m}^{2}$, compared to Optimized Attention. From this and Rem. 3.2, it follows that Efficient Attention has $h+h=2h$ matrix multiplication and $d_{m}^{2}+d_{m}^{2}=2d_{m}^{2}$ parameters fewer than standard attention. ∎